OpenStudy (abbycross167):

Can someone please help me with this math question? Please? I will fan and give a medal! Question- Simplify the following exponential expressions. – (4x^5y)^0 (5xy)^-3

OpenStudy (anonymous):

– (4x^5y)^0 = 1 (5xy)^-3 = 1/(5xy)^3 so 1 * 1/(5xy)^3 = 1/(5xy)^3